OpenStudy (anonymous):

Evaluate the function f(x)=7-x/-5+x at the indicated value f(a)

OpenStudy (anonymous):

If \[f(x)=\frac{ 7-x }{ 5+x }\] Then f(a) = and replace all the x's with a

OpenStudy (anonymous):

i tried but cant seem to get the answer correct ive tried 30 times.

OpenStudy (anonymous):

Did you put \[ f(a)=\frac{ 7-a }{ 5+a }\]? If so, go back and look for a copy error

OpenStudy (anonymous):

sorry, i missed the negative in front of the 5